OpenStudy (anonymous):

What is the value of the expression when a = 5 and b = 3? 4a^2 5b^2

OpenStudy (alexandervonhumboldt2):

substitute 5 for a and 3 for b: 4*5^2*5*3^2=4*25*5*9=simplify

OpenStudy (jacob902):

Substitute a=5 and b=3 into your equation like so: 4*(5)^2 - 5*(3)^2 = 4*25 - 5*9 = 100-45 = 55

OpenStudy (jacob902):

when you see a negative sign in the index (the power) it means you have to invert the expression. So the expression becomes 1/(3^4), because when you invert the negative disappears. Simplifying this it now becomes 1/81, which is approx. 0.0123456789... choose 1/81